OpenStudy (anonymous):

Describe how a gallon of 25% solution is added to an unknown amount of 10% solution to get a 15% solution

OpenStudy (anonymous):

\[ 0.25x+0.10y=0.15(x+y) \]And \[ x=1 \]Solve for \(y\)

OpenStudy (anonymous):

Is that the answer

OpenStudy (anonymous):

It's a method to find the solution.

OpenStudy (anonymous):

What should I right down then